OpenStudy (anonymous):

please help me? How do I simplify these radical expressions? The problem is in the comments, and can someone please explain how to solve it so I know for future reference? Thank you!

OpenStudy (anonymous):

\[\sqrt{144} - \sqrt{40} - \sqrt{10}\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Thank you so much! I appreciate the help!

OpenStudy (anonymous):

welcome.